# Calibrated Physics-Informed Uncertainty Quantification

### v1

@@ H.3 Calibration and Validation
Appendix I Utilising PRE-CP as a measure of model quality
While evaluating the performance of a neural-PDE, it is important to their fit not just to the data but to the underlying physics. PRE-CP will provide guaranteed coverage irregardless of the quality of the model. It will have considerably wider error bounds when the neural-PDE (whether PINN or a Neural Operator) fails to comply with the physics. However, we believe that this is an advantage of our method. In PRE-CP formulation, the bounds are estimated across the PDE residual, where the ground truth for a well-fit solution should always be near zero. If we get wide error bars further away from the 0 for potentially high coverage estimates, it is a strong indication that statistically the model violates the physics of interest.
@@ Appendix I Utilising PRE-CP as a measure of model quality
Figure 18: PRE-CP provides guaranteed coverage irrespective of the model performance, however, the width of the obtained coverage bounds indicates the accuracy of the model in obeying the underlying physics. Coverage taken for $\alpha=0.1\sim 90\%$ coverage.
@@
Figure 22: Navier-Stokes Equations: CP using the Continuity Equation 11 as the PRE.
@@
We train a 2D multi-variable FNO to map the spatio-temporal evolution of the 6 field variables collectively. We deploy an auto-regressive structure that performs time rollouts allowing us to map the initial distribution recursively up until the $20^{th}$ time instance with a step size of 1. Each Fourier layer has 8 modes and a width of 16. The FNO architecture can be found in Table 15. We employ a linear range normalisation scheme, placing the field values between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma & Ba, 2015) with a step-decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using an LP-loss (Gopakumar et al., 2024b). The performance of the trained model can be visualised in Figures 23 and 24.
@@
Figure 23: Navier-Stokes Equations: Temporal evolution of velocity and pressure modelled using the numerical spectral solver (top of the figure) and that of the FNO (bottom of the figure). (Continued on next page)
@@
Figure 24: Navier-Stokes Equations: Temporal evolution of velocity and pressure modelled using the numerical spectral solver (top of the figure) and that of the FNO (bottom of the figure). (Continued from previous page)
@@
For this system, we compute physics residual errors (PRE) for the temperature equation (equation 3 from (Gopakumar et al., 2024b)), as it comprises all the variables modelled by the neural PDE. Empirical coverage obtained by performing PRE-CP over the JOREK FNO is shown in Figure 29.
Figure 29: Empirical coverage obtained by performing PRE-CP over JOREK-FNO
@@
PRE-CP is performed by evaluating the Grad-Shafranov equation over the solution space of the surrogate model, with inputs generated by sampling within the bounds of the PF coil locations. By using the GS equation, we can identify the predictions from the surrogate model that generate untenable equilibria. This allows us to explore the design space quickly while adding trustworthiness to your surrogate model. Both marginal and coverage obtained using PRE-CP with GS equations over the surrogate model is indicated in Figure 34. Marginal-CP shows smooth coverage as it represents the coverage averaged across each cell.
Figure 34: Marginal and joint empirical coverage obtained by performing PRE-CP over the Grad-Shafranov equation

### v2

@@ H.3 Calibration and Validation
Appendix I Utilising CP-PRE as a measure of model quality
While evaluating the performance of a neural-PDE, it is important to their fit not just to the data but to the underlying physics. CP-PRE will provide guaranteed coverage irregardless of the quality of the model. It will have considerably wider error bounds when the neural-PDE (whether PINN or a Neural Operator) fails to comply with the physics. However, we believe that this is an advantage of our method. In CP-PRE formulation, the bounds are estimated across the PDE residual, where the ground truth for a well-fit solution should always be near zero. If we get wide error bars further away from the 0 for potentially high coverage estimates, it is a strong indication that statistically the model violates the physics of interest.
@@ Appendix I Utilising CP-PRE as a measure of model quality
Figure 16: CP-PRE provides guaranteed coverage irrespective of the model performance, however, the width of the obtained coverage bounds indicates the accuracy of the model in obeying the underlying physics. Coverage taken for $\alpha=0.1\sim 90\%$ coverage.
@@
(a) PRE: Momentum Equation
(b) Marginal Bounds
(c) Joint Bounds
@@
Figure 21: Navier-Stokes: CP using the Continuity Equation 12 as the PRE for a neural PDE surrogate model trained to model fluid dynamics. Figure 20(a) depicts the PRE, Figure 20(b) depicts the upper error bar, marginal for each cell, while Figure 20(c) indicates the upper error bar obtained across the entire prediction space. Both are estimated for $90\%$ coverage.
@@
We train a 2D multi-variable FNO to map the spatio-temporal evolution of the 6 field variables collectively. We deploy an auto-regressive structure that performs time rollouts allowing us to map the initial distribution recursively up until the $20^{th}$ time instance with a step size of 1. Each Fourier layer has 8 modes and a width of 16. The FNO architecture can be found in Table 15. We employ a linear range normalisation scheme, placing the field values between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma & Ba, 2015) with a step-decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using an LP-loss (Gopakumar et al., 2024b). The performance of the trained model can be visualised in Figures 22 and 23.
@@
Figure 22: MHD Equations: Temporal evolution of velocity and pressure modelled using the numerical solver (top of the figure) and that of the FNO (bottom of the figure). (Continued on next page)
@@
Figure 23: MHD Equations: Temporal evolution of velocity and pressure modelled using the numerical solver (top of the figure) and that of the FNO (bottom of the figure). (Continued from previous page)
@@
For this system, we compute physics residual errors (PRE) for the temperature equation (equation 3 from (Gopakumar et al., 2024b)), as it comprises all the variables modelled by the neural PDE. Empirical coverage obtained by performing CP-PRE over the JOREK FNO is shown in Figure 29.
Figure 29: Empirical coverage obtained by performing CP-PRE over JOREK-FNO
@@
CP-PRE is performed by evaluating the Grad-Shafranov equation over the solution space of the surrogate model, with inputs generated by sampling within the bounds of the PF coil locations. By using the GS equation, we can identify the predictions from the surrogate model that generate untenable equilibria. This allows us to explore the design space quickly while adding trustworthiness to your surrogate model. Both marginal and coverage obtained using CP-PRE with GS equations over the surrogate model is indicated in Figure 34. Marginal-CP shows smooth coverage as it represents the coverage averaged across each cell.
Figure 34: Marginal and joint empirical coverage obtained by performing CP-PRE over the Grad-Shafranov equation